OpenStudy (anonymous):

A country's population in 1990 was 37 million. In 1999 it was 40 million. Estimate the population in 2016 using the exponential growth formula. Round to the nearest million.

OpenStudy (anonymous):

p=Ae^kt

OpenStudy (anonymous):

so population is: \[P = Ae^{kt}\] The name of the game is to figure out what A and k is. Lets use 1990 as the base year (t = 0). We know that the population is 37 million (so i'll use P = 37 so we dont have zeros flying all over the place, and we just have to remember, P is in millions). Plugging this into the equation we obtain: \[p = Ae^{kt} \Rightarrow 37 = Ae^{k*0} \Rightarrow 37 = Ae^0 \Rightarrow 37 = A\] \[e^0 = 1\]

OpenStudy (anonymous):

So we now know that A = 37 (you should just try to remember that A is always the intial value, so you dont have to repeat the same process over and over)

OpenStudy (anonymous):

Now, to figure out what k is, we use that other set of data. in 1999 there were 40 million people. 1999 is 9 years after 1990 (our base year), so we say t = 9. Remember, P is in millions, so P = 40. So now our equation looks like: \[40 = 37e^{k*9} \Rightarrow \frac{40}{37} = e^{9k} \Rightarrow \ln(\frac{40}{37}) = 9k \Rightarrow \frac{1}{9}\ln(\frac{40}{37}) = k\] I dont have a calculator on me so i cant get that value =/ but thats it. Once you get the number k, you will have your equation. 2016 is 26 years after 1990, so plug in t = 26 and that will be your answer.

OpenStudy (anonymous):

I got a decimal: 0.039

OpenStudy (anonymous):

But those are essentially the same thing, so idk.. Thanks for spelling it out for me though it will help bunches

OpenStudy (anonymous):

what was the answer in the end?

OpenStudy (anonymous):

The answer would be 46 million.